OpenStudy (avanti):

PRECALC QUESTION! Convert the polar representation of this complex number into its standard form: z = 2(cos 11pi/6+ i sin 11pi/6)? A. sqrt(3) -i B. 1- sqrt(3) i C. -sqrt(3)/ 2 -(1/2) i D. -sqrt(3) + i

OpenStudy (psymon):

Find the cos and sin of 11pi/6 and substitute that inside of the parenthesis, then just distribute the 2 inside.

OpenStudy (avanti):

I dont know how.. I'm getting a weird answer

OpenStudy (psymon):

DO you know the values of cos and sin at 11pi/6?

OpenStudy (avanti):

yes cos 11pi/6 is sqrt(3) /2 and sin is (-1/2)

OpenStudy (psymon):

Correct. So then we would substitute that and have. \[2(\frac{ \sqrt{3} }{ 2}-i \frac{ 1 }{ 2 })\] Now just multiply the 2 in :3

OpenStudy (avanti):

Ohh okay got it, thank you!! :D

OpenStudy (psymon):

Yep, not that bad :P

OpenStudy (psymon):

Good luck ^_^